Sung’s Coffee Shop sells a cup of tea for $1.05, a cup of coffee for $1.35, and a cup of hot chocolate for
$2.25. One busy day, 20 more cups of coffee than cups of hot chocolate were sold, and 30 more cups of
coffee than tea, for a total of $202.50 for all three hot drinks. How many cups of each drink were sold?

Found 2 solutions by ewatrrr, josgarithmetic:
Answer by ewatrrr(24785)


Hi
Tea: 1.05
Coffee: 1.35  
Hot Choc: 2.25
 1.05(x-30) + 2.25(x-20) + 1.35x = 202.50
         x = (202.50 + 31.50 + 45)/4.65 = 60 Coffees, 30 Teas & 40 Hot choc
 
Wish You the Best in your Studies.


Answer by josgarithmetic(39616)

Just a way to get started:
DRINK             PRICE         QUANTITY
tea               1.05          h+20-30
coffee            1.35          h+20
hot choc.         2.25          h